BaconLover:

Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a car for a week and driving it x miles is shown below: f(x) = 90 + 0.11x What is the value of f(300), and what does f(300) represent?

BaconLover:

@Warriorz13

1 attachment
Warriorz13:

So I don't think A is a function because 4 square rooted is 2 yes, but 4 square rooted isn't -2. the sqr of 9 is 3, not -3. For B, f(x) = 90 + 0.11x plug in 300 for x f(300) = 90 + 0.11(300) 90 is the cost of the car, 0.11 is the cost per mile so x would be any number he drove, it says to use 300, so it implies he drove 300 miles.

Warriorz13:

A is not a function because 4 square rooted is not -2, and the square root of 9 is 3 not -3. f(x) = 90 + 0.11x plug in 300 for x f(300) = 90 + 0.11(300), 90 would be the cost of the car and the .11 is the cost per mile so the last value, or x would be the number of miles he drove, so the number 300 means he drove 300 miles.